# Mapping Regional Disparities in Discounted Grocery Products

### v1

@@ Abstract
Food waste represents a major challenge to global climate resilience, accounting for almost 10% of annual greenhouse gas emissions. The retail sector is a critical player, mediating product flows between producers and consumers, where supply chain inefficiencies can shape which items are put on sale. Yet how these dynamics vary across geographic contexts remains largely unexplored. Here, we analyze data from Denmark’s largest retail group on near-expiry products put on sale. We uncover the geospatial variations using a dual-clustering approach. We identify multi-scale spatial relationships in retail organization by correlating store clustering – measured using shortest-path distances along the street network – with product clustering based on promotion co-occurrence patterns. Using a bipartite network approach, we identify three regional store clusters, and use percolation thresholds to corroborate the scale of their spatial separation. We find that stores in rural communities put meat and dairy products on sale up to 2.2 times more frequently than metropolitan areas. In contrast, we find that metropolitan and capital regions lean toward convenience products, which have more balanced nutritional profiles but less favorable environmental impacts. By linking geographic context to retail inventory, we provide evidence that reducing food waste requires interventions tailored to local retail dynamics, highlighting the importance of region-specific sustainability strategies.
@@ Introduction
The growing field of food informatics – which draws on large-scale digital records to chart the complex system of human interactions with food products – has transformed our understanding of food processing, consumption, and diet [1, 2, 3, 4]. For example, advances in computational approaches now allow us to predict the levels of food processing [5], revealing that more than 73% of the U.S. food supply is ultra-processed [6], and uncovering the vast, unmapped chemical diversity of our diets [7]. At the same time, integrating mobility data with food choice patterns [8, 9] has shown that the food environments people encounter throughout their daily routines differ markedly from those near their homes, significantly shaping dietary decisions and consumption behaviors [10]. Yet, while this line of research has deepened our understanding of the connections between food and health [11, 12, 13, 14], another dimension of food-related problems emerges as one of the most pressing challenges in modern food systems: food waste, which creates environmental burdens that exacerbate resource scarcity and health risks [15, 16, 17, 18, 19, 20, 21]. Globally, nearly one-third of all food produced for human consumption is lost or discarded each year, with profound environmental, economic, and social consequences [22, 23, 24, 25].
@@
While food waste originates from inefficiencies throughout the supply chain [26, 27, 28, 29] – from production through distribution to consumption – the retail sector represents a critical transition point where products shift from distribution networks to consumers [30, 31, 32]. As a complex adaptive system [33, 34], retail sales are shaped by both promotional interventions on perishable items and evolving consumer preferences [30, 31], creating heterogeneous spatial and temporal patterns in product distribution across stores [35]. For instance, large-scale analyses of grocery purchasing records have linked dietary behavior to health outcomes, revealing spatially structured differences within cities [36], while lifestyle-based consumption behaviors form distinct clusters across urban environments [37], indicating that food choices – and by extension food waste – are deeply embedded in cities’ social and spatial fabric [38, 39]. Given extensive research demonstrating how hierarchical spatial organization [40] fundamentally shapes urban mobility [41, 42], economic activity [43], and overall human behavior patterns [44, 45, 46, 47, 48], we were inspired to question whether or not similar spatial hierarchies also structure the spatial distribution of promotions of perishable foods – in other words, how food waste dynamics and promotional strategies vary across different spatial scales and geographic contexts [39].
@@
We address these questions by analyzing spatial patterns of promotional strategies for near-expiry products across Denmark. Even in a small country like Denmark, the impact of food waste is significant – contributing to approximately 2.3 million tonnes of $\mathrm{CO_{2}}$-equivalent greenhouse gas emissions annually [49], and beyond economic loss and resource depletion, addressing it is essential for achieving sustainable development goals related to climate action, resource efficiency, and social equity [50, 28, 35, 24, 51]. Furthermore, a previous analysis of Danish retail data focusing on alcohol consumption demonstrated significant consumer heterogeneity in food purchasing patterns [52], with distinct lifestyle-based clusters showing different preferences for healthy versus processed foods, highlighting how food patterns could vary across regions with fundamentally different consumer bases. Products are discounted by retailers as part of a complex optimization strategy that balances supply constraints, demand variability, and expiration pressures to minimize waste while maximizing revenue [53, 54, 55, 56]. When a product is placed on sale repeatedly over time, this temporal pattern signals systematic inefficiencies in demand forecasting models, inventory optimization protocols, and product lifecycle management strategies [56]. Consequently, understanding the spatial heterogeneity of promotional patterns across geographic scales offers a diagnostic lens into how place-based variations in logistics, forecasting, and consumer behavior shape food waste trajectories in retail networks.
@@ Results
Salling Group, Denmark’s largest retailer, has focused on food waste reduction with a goal to cut food waste by 50% by 2030 [57]. As part of this strategy, Salling Group provides a public API (Application Programming Interface) that offers real-time information about discounted products approaching their expiration dates, such as meat, dairy products, and prepared meals. We collected data on discounted products across Netto stores, which operate 542 locations in Denmark and rank among the country’s top discount supermarket chains (See Supplementary Materials, Section 1 and Methods for data description). Our dataset captures daily information for each product, recording when it is on sale at a discounted price and where. This dataset allows us to examine the complex spatial-temporal patterns of food waste across Denmark over our 153-day period (See Methods for further details). Furthermore, we manually inspect the dataset and remove offers corresponding to high volumes of products (>50>50 units), representing the 95th percentile, as these tend to correspond to large stock quantities on sale for promotional campaigns rather than genuine near-expiry products. We combine store location data with OpenStreetMap to compute distances between stores, and each product has been mapped through its corresponding Nutri-Score and Environmental impact scores to enable comprehensive sustainability analysis using OpenFoodFacts (See Methods for further details).
@@
To investigate regional product distribution patterns across Denmark, we analyze data spanning the country’s five administrative regions: Hovedstaden (Capital Region), Sjælland (Zealand), Syddanmark (Southern Denmark), Midtjylland (Central Jutland), and Nordjylland (North Jutland) (See Figure 1 a). First, we employ hierarchical spatial clustering of store locations using shortest-path distances on the street network to identify densely connected urban cores and their peripheral hinterlands [58, 40]. We identify groups of stores that are densely connected in terms of street-network travel distance, while leaving more isolated locations unclustered. This approach allows us to distinguish compact urban cores from their (diffuse) peripheries (see Methods for full details). Figure 1 b shows a distinct core–periphery structure [59], as it is characterized by high silhouette scores (0.65), suggesting well-separated and coherent clusters of stores. The maximum persistence (0.36), which quantifies the robustness of clusters across multiple resolutions, is low-to-moderate, reflecting that while the Capital Region forms a stable dominant core, peripheral areas display fragmented, low-persistence clustering that aligns with their scattered population distributions.
To quantify both the magnitude of store-product promotional dynamics – reflecting inventory management decisions and consumer behavior patterns – and the temporal persistence of discounting activities that signal systematic surplus conditions, we aggregate products by category and compute cumulative daily promotion frequencies across the observation period. Figure 1 c shows that Cheese, Pork, Ready-to-Eat Meals, Breads, and Yoghurt emerge as the most commonly offered items nationwide. These categories represent consumer staples with varying shelf lives, ranging from perishable items such as Pork and Bread, which typically remain viable for approximately one week, to longer-lasting products like Cheese and Yoghurt. Although overall patterns of on sale items are similar across regions, normalized entropy highlights regional distinctions: the Capital Region (Hovedstaden) – comprising the urban core of Copenhagen and Frederiksberg alongside adjacent rural municipalities such as Fredensborg, Hillerød, and Gribskov – have a low mean entropy (0.79) and greatest variability. This means that fewer products tend to be on sale, but that those products differ markedly across stores, indicating substantial heterogeneity in promotional diversity within the region. On the other end of the scale Syddanmark (0.820), Midtjylland (0.818), Nordjylland (0.815), and Sjælland (0.811) display higher and more stable diversity. These patterns suggest regional differences with respect to on sale product diversity [33, 53, 54, 55, 56].
@@
Building upon these entropy-based findings, we characterize inter-store differences in promotional composition by computing the Bray-Curtis [60] dissimilarity matrices across all store pairs, using category-specific promotion frequencies as abundance proxies that reflect underlying supply-demand dynamics (See Methods and Supplementary Materials, Section 2). The Bray-Curtis dissimilarity index quantifies compositional differences between stores by comparing the relative frequencies of promotional activities across product categories, in direct analogy to ecological analyses that assess community dissimilarity based on species abundance distributions. Product categories with high promotional frequencies – indicative of persistent surplus conditions or systematic demand forecasting errors – contribute disproportionately to inter-store dissimilarity, with the index ranging from zero (identical promotional profiles) to one (completely distinct promotional patterns), thereby allowing us to understand the spatial heterogeneity of supply chain inefficiencies across retail networks. Then, we assess the relationship between spatial proximity and this similarity using Mantel correlations [61] across multiple distance scales (See Methods), to construct a cumulative correlogram to evaluate whether stores that are geographically closer to one another also tend to have more similar compositions (See Figure 1 d).
@@
We observe a positive spatial-product correlations at smaller scales, peaking at around $4000$ m (Mantel Correlation $\sim 0.1$, $p<0.05$), driven primarily by the cohesive Capital Region cluster (See Figure 1 d I). This correlation diminishes upon inclusion of the Sjælland region (See Figure 1 d II) and becomes zero and nonsignificant ($p>0.05$) when peripheral regional clusters merge with the largest connected component ($\sim 26500$ m). Beyond this critical distance, the correlation starts to be statistically significant, but in the negative direction, indicating divergence in composition among more distant stores (See Figure 1 d III). This scale-dependent pattern suggests that while local retail environments in the capital region exhibit modest product homogeneity, regional differentiation strategies emerge at larger spatial scales, indicating hierarchical organization in Denmark’s retail landscape [40]. This hierarchical organization is statistically confirmed by PERMANOVA analysis ($F=6.9$, $p=0.001$), demonstrating that spatial clusters correspond to genuinely distinct composition (See Methods for further details). To assess the robustness of our findings, we perform a sensitivity analysis by excluding random days from the dataset, incorporating an alternative check that measures total product availability as the sum of individual volumes offered, and normalizing volumes by dividing it by the corresponding population for each store within its respective municipality (See Methods and Supplementary Materials, Section 3). The analysis is unchanged in the light of these perturbations.
@@
Building on the fact that the spatial clustering results reveal scale-dependent patterns in Denmark’s retail geography, we use a bipartite network approach to identify which products drive the regional differences (See Figure 2a). The bipartite network framework avoids circular reasoning by allowing product–store associations to emerge organically from data rather than relying on geographical distance, which risks conflating spatial proximity with functional relationships. The raw network is dense – comprising 542 stores, 1,854 products, and over 418,000 weighted links, where each edge weight represents the total number of days a specific product appears on sale in a given store – many of which connections may be redundant or statistically uninformative. To prune the network we apply two filtering strategies: Revealed Comparative Advantage (RCA) and then the Bipartite Configuration Model (BiCM) to extract the underlying structure from the network [62, 63, 64]. RCA measures whether a store offers products more frequently than expected given the overall market [65], with values >1>1 indicating relative specialization and reflecting local market adaptation with their demographic context (See Methods and Supplementary Materials, Section 4 for the RCA distribution). After removing weaker connections through RCA filtering, we employ BiCM to statistically validate the observed network structure. In a nutshell, BiCM generates an ensemble of randomized bipartite networks that preserve the degree sequences of stores (total number of unique products connected) and products (total number of unique stores connected) while randomizing other structural features (See Methods for further details). This approach constitutes a null model used to distinguish genuine patterns from artifacts of store size or product popularity (defined here as the number of stores in which a product is on sale) [64].
@@
When projecting the validated bipartite structure onto the store layer, we identify three well-defined communities (modularity $Q=0.49$). These clusters correspond closely to the three phases previously revealed as part of our purely spatial analysis: Metropolitan Areas, Capital Region, and Countryside (See Figure 2b). Notably, the Metropolitan Areas cluster extends beyond Copenhagen, encompassing cities such as Odense and Aarhus [59], while the Capital Region cluster reflects the Finger Plan [66], with urban development radiating from central Copenhagen along five suburban ‘fingers’ separated by green wedges. These communities are robust and emerge from two different community detection methods, Louvain and Infomap [67, 68] (Adjusted Rand Index $ARI=0.82$, $p<0.001$). We incorporate distance weights into the store–store network and apply percolation analysis [40] to identify the hierarchical connectivity structure of this network. This percolation profile captures how densely connected store clusters emerge as we vary the distance threshold parameter. Notably, this distance-weighted percolation profile again closely mirrors that of our purely spatial percolation analysis, indicating that spatial proximity information is inherently encoded within the product-based similarity network (See Figure 2c). Collectively, this convergence reveals that spatial proximity and product on sale similarity offer complementary perspectives on the same underlying organizational principle: stores operating within similar geographical contexts face comparable market demands and operational constraints, which in turn generate convergent promotional strategies that reflect shared supply chain challenges and consumer preference patterns [33, 37, 39]. As in the preceding section, we perform a sensitivity analysis on the bipartite network weights, as detailed in Supplementary Materials, Section 4, which further underscores the robustness of our results.
@@
On the other hand, the product layer projection identifies co-promotional clusters – groups of items that systematically appear on sale together across multiple stores – thereby revealing regionally-characteristic product communities and identifying critical “bridge” products that link metropolitan, suburban, and rural retail environments [34]. This projection has a very high modularity ($Q=0.83$) and again we find agreement between Infomap and Louvain community assignments (Adjusted Rand Index $ARI=0.68$, $p<0.001$). The contrast between store-layer modularity ($Q=0.49$) and product-layer modularity ($Q=0.83$) points to moderate asymmetric organizational forces [63]: while geographical constraints create moderate store clustering, product relationships exhibit much stronger functional cohesion. This cohesion implies that promotional co-occurrence patterns – driven by shared supply chain dynamics and consumer purchasing behaviors – create more rigid boundaries in product networks than spatial proximity generates in store networks. Although more than one hundred distinct communities emerge, this (large) number mainly reflects the fragmented structure of the network (See Figure 2 d). Many communities align closely with conventional categories – Light and Rye Breads, Ready-to-Eat Meals, Beverages, and various Meat assortments – while notable divergences appear, such as a central cluster that unites Yoghurt, Cheese, Milk, Tapas and Whipped Cream. Focusing on the largest connected component, which comprises 433 products, we compute node betweenness centrality (bc) to identify core “hub” products that frequently are part of shortest paths between items. In contrast to the categories ranked in Figure 1 c, which reflect aggregate promotional volume (total frequency of discounting events), here we identify structural hub products that bridge diverse promotional clusters and orchestrate both regional homogeneity and inter-regional differentiation through their network connectivity patterns. Categories such as Dressings, Meat, Pâté, and Breads are most likely to include nodes with high betweenness centrality ($\text{bc}>0.03$, top 90th percentile; see Figure 2e), suggesting they serve as key intermediaries.
@@
Building on the identification of store communities and product clusters, we now turn to the question of what differentiates these clusters in practice [63]. To address this, we evaluate the per-capita on sale frequency, which reflects how frequently product groups (product categories, nutritional profiles, or environmental ratings) go on sale relative to the number of customers each store serves. The per-capita on sale frequency is defined as the ratio between validated store–product connections within each community $c$ and product category $k$ (or Nutri-score and Environmental Score), $E_{ck}$, and the total population $P_{c}$ of distinct stores in the community $c$. We scale this index by a factor of 100,000, so that it reflects the number of on sale items per 100,000 potential customers. The per-capita on sale frequency quantifies both systematic oversupply conditions (indicating demand forecasting failures and waste risk) and regionally differentiated availability patterns that mirror localized consumer preferences. Thus this measure is able to capture the dynamic relationship between supply chain strategies and community characteristics.
@@
Figure 3 a shows systematic differences in promotional patterns across the three community types, each exhibiting distinct profiles that reflect underlying supply-demand dynamics and regional consumption preferences [52, 37]. A Countryside versus Metropolitan Areas comparison reveals systematic differences in the promotional intensity of meat: Countryside shows a total meat frequency of 0.47 – 2.2 times higher than the Metropolitan Areas’ 0.21. In fact Countryside dominates across all meat categories: Pork is nearly twice as frequent as in Metropolitan Areas, while Chicken is absent in the Metropolitan Area. These patterns suggest that the Countryside has higher promotional intensities for specific meat categories – Pork, Chicken, Lamb, and Sausage – than Metropolitan Areas. Additionally, Butter shows markedly higher promotional intensity in the Countryside, with per-capita on sale frequencies two times greater than those observed in the Metropolitan Areas. Conversely, products such as Tapas, Pasta, and Crackers are more frequently observed in the Capital Region and the Metropolitan Areas. Additionally, the Capital Region demonstrates a pronounced preference for convenience and on-the-go products: Cold Coffee, Dessert Snacks, and Crackers all rank among its top five items.

### v2

@@ Abstract
Food waste represents a major challenge to global climate resilience, accounting for almost 10% of annual greenhouse gas emissions. The retail sector is a critical player, mediating product flows between producers and consumers, where supply chain inefficiencies can shape which items are put on sale. Yet how these dynamics vary across geographic contexts remains largely unexplored. Here, we analyze data from Denmark’s largest retail group on near-expiry products put on sale. We uncover the geospatial variations using a dual-clustering approach. We characterize multi-scale spatial relationships in retail organization by correlating store clustering – measured using shortest-path distances along the street network – with product clustering based on promotion co-occurrence patterns. Using a bipartite network approach, we identify three regional store clusters, and use percolation thresholds to corroborate the scale of their spatial separation. We find that stores in rural communities put meat and dairy products on sale up to 2.2 times more frequently than metropolitan areas. In contrast, metropolitan and capital regions lean toward convenience products, which have more balanced nutritional profiles but less favorable environmental impacts. By linking geographic context to retail inventory, we provide evidence that reducing food waste requires interventions tailored to local retail dynamics, highlighting the importance of region-specific sustainability strategies.
@@ Introduction
The growing field of food informatics – which draws on large-scale digital records to chart the complex system of human interactions with food products – has transformed our understanding of food processing, consumption, and diet [tufano_data-driven_2025, tilman_global_2014, ispirova_informatics_2025, kuhl_ai_2025]. For example, advances in computational approaches now allow us to predict the levels of food processing [menichetti_machine_2023], revealing that more than 73% of the U.S. food supply is ultra-processed [ravandi_prevalence_2025], and uncovering the vast, unmapped chemical diversity of our diets [barabasi_unmapped_2019]. At the same time, integrating mobility data with food choice patterns [horn_population_2023, xu_integrating_2023] has shown that the food environments people encounter throughout their daily routines differ markedly from those near their homes, significantly shaping dietary decisions and consumption behaviors [garcia_bulle_bueno_effect_2024]. Yet, while this line of research has clarified the link between food and health through nutrition [wingrove_outcomes_2025, lane_ultra-processed_2024, muncke_health_2025, fanzo_sustainable_2022], food systems also influence human health through another key pathway: food waste, which generates environmental burdens that exacerbate resource scarcity and associated health risks [minor_rising_2022, romanello_2024_2024, nijdam_price_2012, springmann_analysis_2016, wang_food_2025, anshassi_improving_2025, shepon_opportunity_2018]. Globally, nearly one-third of all food produced for human consumption is lost or discarded each year, with profound environmental, economic, and social consequences [macdiarmid_food_2022, ambikapathi_global_2022, schneider_state_2023, crippa_food_2021].
@@
While food waste originates from inefficiencies throughout the supply chain [nerin_food_2016, huang_food_2021, campi_specialization_2021, filimonau_exploratory_2017] – from production through distribution to consumption – the retail sector represents a critical transition point where products shift from distribution networks to consumers [principato_food_2018, hubner_demand_2013, tian_extracting_2021]. As a complex adaptive system [fox_consumer_2004, pennacchioli_retail_2014], retail sales are shaped by both promotional interventions on perishable items and evolving consumer preferences [principato_food_2018, hubner_demand_2013], creating heterogeneous spatial and temporal patterns in product distribution across stores [moragues-faus_urban_2021]. For instance, large-scale analyses of grocery purchasing records have linked dietary behavior to health outcomes, revealing spatially structured differences within cities [aiello_large-scale_2019], while lifestyle-based consumption behaviors form distinct clusters across urban environments [lee_we_2025], indicating that food choices – and by extension food waste – are deeply embedded in cities’ social and spatial fabric [dong_social_2018, seto_hidden_2016]. Given extensive research demonstrating how hierarchical spatial organization [arcaute_cities_2016] fundamentally shapes urban mobility [lenormand_comparing_2015, bassolas_hierarchical_2019], economic activity [di_clemente_urbanization_2021], and overall human behavior patterns [lengyel_role_2020, yang_identifying_2023, pan_urban_2013, bettencourt_urban_2010, sapienza_exposure_2023], we were inspired to question whether or not similar spatial hierarchies also structure the spatial distribution of promotions of perishable foods – in other words, how food waste dynamics and promotional strategies vary across different spatial scales and geographic contexts [seto_hidden_2016].
@@
We address these questions by analyzing spatial patterns of promotional strategies for near-expiry products across Denmark. Even in a small country like Denmark, the impact of food waste is significant – contributing to approximately 2.3 million tonnes of $\mathrm{CO_{2}}$-equivalent greenhouse gas emissions annually [pacecircular_denmark_foodwaste], and beyond economic loss and resource depletion, addressing it is essential for achieving sustainable development goals related to climate action, resource efficiency, and social equity [otto_social_2020, campi_specialization_2021, moragues-faus_urban_2021, schneider_state_2023, schneeman_national_2020]. Furthermore, a previous analysis of Danish retail data focusing on alcohol consumption demonstrated significant consumer heterogeneity in food purchasing patterns [johansen_food_2006], with distinct lifestyle-based clusters showing different preferences for healthy versus processed foods, highlighting how food patterns could vary across regions with fundamentally different consumer bases. Products are discounted by retailers as part of a complex optimization strategy that balances supply constraints, demand variability, and expiration pressures to minimize waste while maximizing revenue [buisman_discounting_2019, kayikci_data-driven_2022, atan_displaying_2025, anderson_how_2019]. When a product is placed on sale repeatedly over time, this temporal pattern signals systematic inefficiencies in demand forecasting models, inventory optimization protocols, and product lifecycle management strategies [anderson_how_2019]. Consequently, understanding the spatial heterogeneity of promotional patterns across geographic scales offers a diagnostic lens into how place-based variations in logistics, forecasting, and consumer behavior shape food waste trajectories in retail networks.
@@ Results
Salling Group, Denmark’s largest retailer, has focused on food waste reduction with a goal to cut food waste by 50% by 2030 [sallinggroup_foodwaste]. As part of this strategy, Salling Group provides a public API (Application Programming Interface) that offers real-time information about discounted products approaching their expiration dates, such as meat, dairy products, and prepared meals. We collected data on discounted products across Netto stores, which operate 542 locations in Denmark and rank among the country’s top discount supermarket chains (See Supplementary Materials, Section 1 and Methods for data description). Our dataset captures daily information for each product, recording when it is on sale at a discounted price and where. This dataset allows us to examine the complex spatial-temporal patterns of food waste across Denmark over our 153-day period (See Methods for further details). Furthermore, we manually inspect the dataset and remove offers corresponding to high volumes of products (>50>50 units), representing the 95th percentile, as these tend to correspond to large stock quantities on sale for promotional campaigns rather than genuine near-expiry products. We combine store location data with OpenStreetMap to compute distances between stores, and each product has been mapped through its corresponding Nutri-Score and Environmental impact scores to enable comprehensive sustainability analysis using OpenFoodFacts (See Methods for further details).
@@
To investigate product distribution patterns across Denmark, we analyze data spanning the country’s five administrative regions: Hovedstaden (Capital Region), Sjælland (Zealand), Syddanmark (Southern Denmark), Midtjylland (Central Jutland), and Nordjylland (North Jutland) (See Figure 1 a). First, we employ hierarchical spatial clustering of store locations using shortest-path distances on the street network to identify densely connected urban cores and their peripheral hinterlands [hutchison_density-based_2013, arcaute_cities_2016]. We identify groups of stores that are densely connected in terms of street-network travel distance, while leaving more isolated locations unclustered. This approach allows us to distinguish compact urban cores from their (diffuse) peripheries (see Methods for full details). Figure 1 b shows a distinct core–periphery structure [marin_scalar_2024], as it is characterized by high silhouette scores (0.65), suggesting well-separated and coherent clusters of stores. The maximum persistence (0.36), which quantifies the robustness of clusters across multiple resolutions, is low-to-moderate, reflecting that while the Capital Region forms a stable dominant core, peripheral areas display fragmented, low-persistence clustering that aligns with their scattered population distributions.
To quantify both the magnitude of store-product promotional dynamics – reflecting inventory management decisions and consumer behavior patterns – and the temporal persistence of discounting activities that signal systematic surplus conditions, we aggregate products by category and compute cumulative daily promotion frequencies across the observation period. Figure 1 c shows that Cheese, Pork, Ready-to-Eat Meals, Breads, and Yoghurt emerge as the most commonly offered items nationwide. These categories represent consumer staples with varying shelf lives, ranging from perishable items such as Pork and Bread, which typically remain viable for approximately one week, to longer-lasting products like Cheese and Yoghurt. Although overall patterns of on sale items are similar across regions, normalized entropy – a measure of diversity where higher values indicate a more uniform distribution of promotional activity across product categories – highlights regional distinctions: the Capital Region (Hovedstaden) – comprising the urban core of Copenhagen and Frederiksberg alongside adjacent rural municipalities such as Fredensborg, Hillerød, and Gribskov – have a low mean entropy (0.79) and greatest variability across stores within the region. This means that, on average, fewer products tend to be on sale, and the high variability suggests substantial heterogeneity in promotional strategies within the region. On the other end of the scale Syddanmark (0.820), Midtjylland (0.818), Nordjylland (0.815), and Sjælland (0.811) display higher and more stable diversity. These patterns suggest regional differences with respect to on sale product diversity [fox_consumer_2004, buisman_discounting_2019, kayikci_data-driven_2022, atan_displaying_2025, anderson_how_2019].
@@
Building upon these entropy-based findings, we characterize inter-store differences in promotional composition by computing the Bray-Curtis [bray_ordination_1957] dissimilarity matrices across all store pairs, using category-specific promotion frequencies as abundance proxies that reflect underlying supply-demand dynamics (See Methods and Supplementary Materials, Section 2). The Bray-Curtis dissimilarity index quantifies compositional differences between stores by comparing the relative frequencies of promotional activities across product categories, in direct analogy to ecological analyses that assess community dissimilarity based on species abundance distributions. Product categories with high promotional frequencies – indicative of persistent surplus conditions or systematic demand forecasting errors – contribute disproportionately to inter-store dissimilarity, with the index ranging from zero (identical promotional profiles) to one (completely distinct promotional patterns), thereby allowing us to understand the spatial heterogeneity of supply chain inefficiencies across retail networks. Then, we examine the relationship between geographic proximity and compositional similarity using Mantel correlations [diniz-filho_mantel_2013] computed across increasing distance thresholds (see Methods). These correlations are summarized in a cumulative correlogram that characterizes how promotional composition similarity varies with spatial separation (Figure 1d).
@@
We observe a positive spatial-product correlations at smaller scales, peaking at around $4000$ m (Mantel Correlation $\sim 0.1$, $p<0.05$), driven primarily by the cohesive Capital Region cluster (See Figure 1 d I). This correlation diminishes upon inclusion of the Sjælland region (See Figure 1 d II) and becomes zero and nonsignificant ($p>0.05$) when peripheral regional clusters merge with the largest connected component ($\sim 26500$ m). Beyond this critical distance, the correlation starts to be statistically significant, but in the negative direction, indicating divergence in composition among more distant stores (See Figure 1 d III). This scale-dependent pattern suggests that while local retail environments in the capital region exhibit modest product homogeneity, regional differentiation strategies emerge at larger spatial scales, indicating hierarchical organization in Denmark’s retail landscape [arcaute_cities_2016]. This hierarchical organization is statistically confirmed by PERMANOVA analysis ($F=6.9$, $p=0.001$), which tests for differences in multivariate composition using permutation-based comparisons of Bray–Curtis dissimilarities, demonstrating that spatial clusters correspond to genuinely distinct product compositions (see Methods for further details). To assess the robustness of our findings, we perform a sensitivity analysis by excluding random days from the dataset, incorporating an alternative check that measures total product availability as the sum of individual volumes offered, and normalizing volumes by dividing it by the corresponding population for each store within its respective municipality (See Methods and Supplementary Materials, Section 3). The analysis is unchanged in the light of these perturbations.
@@
Building on the fact that the spatial clustering results reveal scale-dependent patterns in Denmark’s retail geography, we use a bipartite network approach to identify which products drive the regional differences (See Figure 2a). The bipartite network framework avoids circular reasoning by allowing product–store associations to emerge organically from data rather than relying on geographical distance, which risks conflating spatial proximity with functional relationships. The raw network is dense – comprising 542 stores, 1,854 products, and over 418,000 weighted links, where each edge weight represents the total number of days a specific product appears on sale in a given store – many of which connections may be redundant or statistically uninformative. To prune the network we apply two filtering strategies: Revealed Comparative Advantage (RCA) and then the Bipartite Configuration Model (BiCM) to extract the underlying structure from the network [saracco_randomizing_2015, saracco_inferring_2017, cimini_statistical_2019]. RCA measures whether a store offers products more frequently than expected given the overall market [balassa_revealed_1989], with values above $1$ indicating relative specialization (See Methods and Supplementary Materials, Section 4 for the RCA distribution). After removing weaker connections through RCA filtering, we employ BiCM to statistically validate the observed network structure. In a nutshell, BiCM generates an ensemble of randomized bipartite networks that preserve the degree sequences of stores (total number of unique products connected) and products (total number of unique stores connected) while randomizing other structural features (See Methods for further details). This approach constitutes a null model used to distinguish genuine patterns from artifacts of store size or product popularity (defined here as the number of stores in which a product is on sale) [cimini_statistical_2019].
@@
When projecting the validated bipartite structure onto the store layer, we identify three well-defined communities (modularity $Q=0.49$, indicating strong internal connectivity within clusters compared to random connections). These clusters correspond closely to the three phases previously revealed as part of our purely spatial analysis: Metropolitan Areas, Capital Region, and Countryside (See Figure 2b). Notably, the Metropolitan Areas cluster extends beyond Copenhagen, encompassing cities such as Odense and Aarhus [marin_scalar_2024], while the Capital Region cluster reflects the Finger Plan [knowles_transit_2012], with urban development radiating from central Copenhagen along five suburban ‘fingers’ separated by green wedges. These communities are robust and emerge from two different community detection methods, Louvain and Infomap [blondel_fast_2008, rosvall_maps_2008] (Adjusted Rand Index $ARI=0.82$, $p<0.001$, See Methods for further details). We incorporate distance weights into the store–store network and apply percolation analysis [arcaute_cities_2016] to identify the hierarchical connectivity structure of this network. This percolation profile captures how densely connected store clusters emerge as we vary the distance threshold parameter. Notably, this distance-weighted percolation profile closely mirrors that of our purely spatial percolation analysis, indicating that spatial proximity information is inherently encoded within the product-based similarity network (See Figure 2c). Collectively, the convergence reveals that a) spatial proximity and b) product on sale similarity offer complementary perspectives on the same underlying organizational principle: stores that operate within similar geographical contexts face comparable market demands and operational constraints. These, in turn, generate convergent promotional patterns. The convergence also highlights potential shared inefficiencies that point to a sharp contrast between rural and urban supply-demand dynamics, and reflecting shared supply chain challenges and consumer preference behaviors [fox_consumer_2004, lee_we_2025, seto_hidden_2016]. For instance, the consistent discounting of meat in rural areas versus convenience foods in the capital reflect distinct local inefficiencies. These disparities suggest that a path to sustainable waste reduction can lie in shifting from reactive uniform discounting to proactive, region-specific inventory adjustments [halloran_addressing_2014]. As in the preceding section, we perform a sensitivity analysis on the bipartite network weights, as detailed in Supplementary Materials, Section 4, which further underscores the robustness of our results.
@@
On the other hand, the product layer projection captures similarities in aggregate promotional patterns across stores, revealing regionally characteristic product communities and identifying critical “bridge” products [pennacchioli_retail_2014]. This projection has a very high modularity ($Q=0.83$) and again we find agreement between Infomap and Louvain community assignments (Adjusted Rand Index $ARI=0.68$, $p<0.001$). The contrast between store-layer modularity ($Q=0.49$) and product-layer modularity ($Q=0.83$) points to moderate asymmetric organizational forces [saracco_inferring_2017]: while geographical constraints create moderate store clustering, product relationships exhibit much stronger functional cohesion. This cohesion implies that promotional co-occurrence patterns – driven by shared supply chain dynamics and consumer purchasing behaviors – create more rigid boundaries in product networks than spatial proximity generates in store networks. Although more than one hundred distinct communities emerge, this (large) number mainly reflects the fragmented structure of the network (See Figure 2 d). Many communities align closely with conventional categories – Light and Rye Breads, Ready-to-Eat Meals, Beverages, and various Meat assortments – while notable divergences appear, such as a central cluster that unites Yoghurt, Cheese, Milk, Tapas and Whipped Cream. Focusing on the largest connected component, which comprises 433 products, we compute node betweenness centrality (bc) to identify core “hub” products that are frequently part of shortest paths between items. Unlike the categories in Figure 1c, which are ranked by overall promotional volume, here we identify structural hub products that connect diverse promotional clusters. These clusters shape both regional uniformity and inter-regional differences through their network connectivity. Categories such as Dressings, Meat, Pâté, and Breads are most likely to include nodes with high betweenness centrality ($\text{bc}>0.03$, top 90th percentile; see Figure 2e), suggesting they serve as key intermediaries.
@@
Building on the identification of store communities and product clusters, we now turn to the question of what differentiates these clusters in practice [saracco_inferring_2017]. To address this, we evaluate the per-capita on sale frequency, which reflects how frequently product groups (product categories, nutritional profiles, or environmental ratings) go on sale relative to the number of customers each store serves. The per-capita on sale frequency is defined as the ratio between validated store–product connections within each community $c$ and product category $k$ (or Nutri-score and Environmental Score), $E_{ck}$, and the total population $P_{c}$ of distinct stores in the community $c$. We scale this index by a factor of 100,000, so that it reflects the number of on sale items per 100,000 potential customers. The per-capita on sale frequency quantifies both oversupply conditions (indicating demand forecasting failures and waste risk) and availability patterns that mirror consumer preferences. Thus this measure is able to capture the relationship between supply chain strategies and community characteristics.
@@
Figure 3 a shows systematic differences in promotional patterns across the three community types, each exhibiting distinct profiles [johansen_food_2006, lee_we_2025]. A Countryside versus Metropolitan Areas comparison reveals systematic differences in the promotional intensity of meat: Countryside shows a total meat frequency of 0.47 – 2.2 times higher than the Metropolitan Areas’ 0.21. In fact Countryside dominates across all meat categories: Pork is nearly twice as frequent as in Metropolitan Areas, while Chicken is absent in the Metropolitan Area. These patterns suggest that the Countryside has higher promotional intensities for specific meat categories – Pork, Chicken, Lamb, and Sausage – than Metropolitan Areas. Additionally, Butter shows markedly higher promotional intensity in the Countryside, with per-capita on sale frequencies two times greater than those observed in the Metropolitan Areas. Conversely, products such as Tapas, Pasta, and Crackers are more frequently observed in the Capital Region and the Metropolitan Areas. Additionally, the Capital Region demonstrates a pronounced preference for convenience and on-the-go products: Cold Coffee, Dessert Snacks, and Crackers all rank among its top five items.